OpenStudy (anonymous):

two integers differ by 12 and the sum of their squares is 74. find the integers two integers differ by 12 and the sum of their squares is 74. find the integers @Mathematics

OpenStudy (anonymous):

5 and 7

OpenStudy (anonymous):

how did you do that though?

OpenStudy (anonymous):

one has to be negative, so -5 and 7 or -7 and 5.

OpenStudy (anonymous):

x-y=12 and x^2+y^2=74, use the first equation and rearrange for x, square it and plug it into the second equation and solve for y. When you have your values for y, plug them into the first equation to get the corresponding x-values

OpenStudy (anonymous):

ok i got x= y+12 but now i dont get it